OpenStudy (anonymous):

6(n+5)=66 6n+5=66 What do I do after that?

jimthompson5910 (jim_thompson5910):

you distributed incorrectly

jimthompson5910 (jim_thompson5910):

6(n+5) = 6*n + 6*5 = 6n + 30

jimthompson5910 (jim_thompson5910):

so your second step should be 6n+30 = 66

jimthompson5910 (jim_thompson5910):

after that Subtract 30 from both sides. then divide both sides by 6 to isolate n.

OpenStudy (anonymous):

N=6??

jimthompson5910 (jim_thompson5910):

n = 6 is correct. Nice job